# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

### v2

@@ Appendix A Adapting FairPO for Multi-Attribute Generation
This section outlines a key direction for future research: adapting the FairPO framework to multi-attribute generation tasks. The methodologies described here are conceptual and represent a planned extension of our work, building directly upon the principles established in the main paper for classification. Experimental validation of this approach is part of our future agenda.
In such a generative setting, the goal would be to produce an output sequence $y$ from a prompt $x$ using a single generative policy $\pi_{\mathbf{w}}(y|x)$, such that the generation aligns with fairness goals defined over a set of attributes $\mathcal{A}$. As in the classification setting, a potential approach involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets. The core GRPO minimax structure (Eq. 18) could be retained, with group losses defined based on a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$ and the attributes $j_{i}$ driving the preferences.
A primary consideration in this adaptation would be how preference losses are applied, particularly for the privileged group, as DPO would then operate on the log-probabilities of entire generated sequences rather than individual label scores.
Proposed Privileged Loss ($\mathcal{L}_{\mathcal{P}}$): For future generative applications, the aim for privileged attributes $j\in\mathcal{P}$ would be to ensure that the learned policy $\pi_{\mathbf{w}}$ strongly reflects observed preferences $y_{w}\succ y_{l}$ when the preference was established based on such an attribute. This could translate to assigning significantly higher relative log-probability to the preferred sequence $y_{w}$ compared to the dispreferred sequence $y_{l}$, relative to a reference policy $\pi_{\text{ref}}$.
This could be achieved by utilizing a standard DPO loss formulation, averaged over the subset of the preference data $\mathcal{D}_{pref}$ where the driving attribute $j$ belongs to the privileged set $\mathcal{P}$. The privileged loss could then be defined as:
@@
Minimizing such a loss would directly encourage the model $\pi_{\mathbf{w}}$ to favor preferred sequences over dispreferred ones for preferences driven by privileged attributes, relative to the reference model $\pi_{\text{ref}}$.
Proposed Non-Privileged Loss ($\mathcal{L}_{\bar{\mathcal{P}}}$): The objective for non-privileged attributes $k\in\bar{\mathcal{P}}$ would remain analogous to the classification setting: to prevent significant degradation compared to a baseline. The non-privileged loss could then utilize a hinge formulation based on the DPO loss for preferences driven by attributes $k\in\bar{\mathcal{P}}$:
@@
This would penalize the model only if its preference modeling for non-privileged attributes degrades substantially beyond the baseline performance (represented by $\log 2$ for random preference) plus a slack $\epsilon^{\prime}$.
The overall FairPO objective for generation would then employ the GRPO minimax formulation to balance $\mathcal{L}_{\mathcal{P}}$ (Eq. 19) and $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq. 20), aiming for robust alignment across both privileged and non-privileged attributes identified in the preference data.
Planned Task and Potential Datasets: Future work will focus on evaluating FairPO for generating text $y$ based on a prompt $x$, aiming for alignment with attributes $\mathcal{A}=\mathcal{P}\cup\bar{\mathcal{P}}$. This would require preference datasets $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}$ where the driving attribute $j_{i}$ for the preference is known. Potential datasets for such an investigation include:
  • Publicly available RLHF datasets (e.g., subsets of Anthropic HH-RLHF [Bai et al., 2022], Summarization preferences [Stiennon et al., 2020]), where attributes like ‘Helpfulness’, ‘Harmlessness’, ‘Conciseness’, ‘Faithfulness’ would need to be partitioned into $\mathcal{P}$ and $\bar{\mathcal{P}}$.
  • Synthetic or newly annotated preference data where attribute labels are explicitly assigned during collection, tailored for this specific fairness investigation.
Prospective Model and Implementation Strategy: We plan to use pretrained generative language models (e.g., Gemma [Team et al., 2024], Llama [Touvron et al., 2023], Mistral [Jiang et al., 2023]) as the base policy $\pi_{\theta}$. These models would then be fine-tuned using the FairPO objective proposed for generation (Equations 19 and 20). The reference policy $\pi_{\text{ref}}$ would typically be the SFT version of the base model.
Proposed Evaluation and Baselines: Evaluation in this future work would primarily rely on preference-based metrics. Comparisons would include:
  • Win rates (human or model-based, e.g., using GPT-4) for preferences driven by privileged attributes $j\in\mathcal{P}$.
  • Win rates for preferences driven by non-privileged attributes $k\in\bar{\mathcal{P}}$.
  • The worst-group win rate across the two attribute groups.
Baselines for comparison would likely include standard non-robust DPO fine-tuning and the SFT reference model.
Future experiments within this generative setting would also investigate the impact of hyperparameters like $\beta$ and $\epsilon^{\prime}$, and the potential effectiveness of adapting the alternative SimPO and CPO loss formulations to the FairPO framework for generation.
@@ Appendix B FairPO Algorithm
The FairPO framework is trained iteratively to solve the minimax objective presented in Eq. 18. The detailed procedure, inspired by the DPO-based variant of FairPO, is provided in Algorithm 2.
The core training loop proceeds as follows:
  1. 1.
@@
    Initialization: Model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}$ are initialized (e.g., from a reference model $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$), and the adaptive group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are typically set to uniform values (e.g., 0.5 each).
  2. 2.
    Iterative Updates: For each training step:
    • An instance $(x_{i},y_{i})$ is sampled from the dataset $\mathcal{D}$. A label $r\in\mathcal{T}$ is then randomly selected for processing from this instance.
    • Privileged Label Processing (if $r\in\mathcal{P}$): The algorithm first identifies if any confusing examples $S_{il}$ exist for the chosen privileged label $l=r$ with respect to other labels (Lines 11-17 in Alg. 2). If $S_{il}\neq\emptyset$, a DPO-inspired preference loss (Eq. 8 or 9) is computed between label $l$ and a sampled confusing example $k\in S_{il}$ (Lines 19-27). This loss encourages the model to correctly rank $l$ against $k$. If $S_{il}=\emptyset$ (no confusing examples found), a standard base classification loss (e.g., BCE, Eq. 10) is used for label $l$ to ensure continued learning (Lines 29-31). The resulting loss contributes to $\mathcal{L}_{\mathcal{P}}^{(s)}$.
    • Non-Privileged Label Processing (if $r\in\bar{\mathcal{P}}$): For a chosen non-privileged label $j=r$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed as per Eq. 17, penalizing deviations from the reference model’s performance beyond a slack $\epsilon$ (Lines 34-37).
@@
    • Adaptive Weight Update (GRPO): The group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated using a mirror ascent step (exponential weighting based on the current step’s group losses $\mathcal{L}_{\mathcal{P}}^{(s)}$ and $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ should be scaled for stable training as explained in Section 4) and then normalized to sum to one (Lines 39-41). This dynamically adjusts the focus on the group currently performing worse.
    • Model Parameter Update: The model parameters $\mathbf{w}_{t}$ for all labels $t\in\mathcal{T}$ are updated via a mirror descent step (gradient descent), using a combined gradient weighted by the updated $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ (Line 42).
  3. 3.
    The process repeats for a predefined number of iterations or until convergence.
@@
This iterative procedure allows FairPO to dynamically balance the specialized objectives for privileged labels against the performance constraints for non-privileged labels, guided by the GRPO mechanism to achieve robust fairness. For variants like FairPO-SimPO or FairPO-CPO, the DPO-inspired preference loss component (Lines 21-26) for privileged labels is replaced with their respective preference formulations (e.g., Eq. 12 or 13, possibly combined with an NLL-like term as in Eq. 14), while the overall GRPO structure and non-privileged handling remain consistent.
@@
1:Initialize: model parameters $\{\mathbf{w}_{t}^{(0)}\in\mathbb{R}^{d}|\forall t\in\mathcal{T}\}$ (e.g., copy $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$), group weights $\alpha_{\mathcal{P}}^{(0)}\leftarrow 0.5,\alpha_{\bar{\mathcal{P}}}^{(0)}\leftarrow 0.5$.
2:Choose: learning rates $\eta_{\mathbf{w}},\eta_{\alpha}$, DPO hyperparameter $\beta$, reference parameters $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$, constraint slack $\epsilon$.
@@
10:          $l\leftarrow r$
11:          $S_{il}^{\text{neg}}\leftarrow\emptyset$, $S_{il}^{\text{pos}}\leftarrow\emptyset$
12:          if $y_{il}=+1$ then $\triangleright$ True Positive case for privileged label $l$
13:              $S_{il}^{\text{neg}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\geq m(x_{i};\mathbf{w}_{l}^{(s)})\}$
14:              $S_{il}\leftarrow S_{il}^{\text{neg}}$
@@
15:          else if $y_{il}=0$ then $\triangleright$ True Negative case for privileged label $l$
16:              $S_{il}^{\text{pos}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\leq m(x_{i};\mathbf{w}_{l}^{(s)})\}$
17:              $S_{il}\leftarrow S_{il}^{\text{pos}}$
18:          end if
@@
19:          if $S_{il}\neq\emptyset$ then $\triangleright$ Confusing examples exist, use DPO-inspired loss
20:              Sample $k\in S_{il}\sim Uniform(\frac{1}{|S_{il}|})$
21:              if $y_{il}=+1$ then $\triangleright$ DPO for True Positive $l$ vs Confusing Negative $k$
22:                  $h_{\mathbf{w}^{(s)}}(x_{i},l,k)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)$.
@@
23:                  $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},l,k)\right)$
24:              else if $y_{il}=0$ then $\triangleright$ DPO for True Negative $l$ vs Confusing Positive $k$
25:                  $h_{\mathbf{w}^{(s)}}(x_{i},k,l)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)$.
@@
26:                  $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},k,l)\right)$
27:              end if
28:              $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{pref}}$
@@
29:              $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{pref}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
30:          else $\triangleright$ No confusing examples, use BCE loss for privileged label $l$
31:              $\mathcal{L}_{\text{BCE}}\leftarrow-[y_{il}\log m(x_{i};\mathbf{w}_{l}^{(s)})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}^{(s)}))]$
32:              $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{BCE}}$
33:              $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{BCE}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
34:          end if
35:      else if $r\in\bar{\mathcal{P}}$ then $\triangleright$ Handle non-privileged label
36:          $j\leftarrow r$
37:          $\ell(\mathbf{w}_{j}^{(s)})\leftarrow-[y_{ij}\log(m(x_{i};\mathbf{w}_{j}^{(s)}))+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}^{(s)}))]$
38:          $\ell(\hat{\mathbf{w}}_{j})\leftarrow-[y_{ij}\log(m(x_{i};\hat{\mathbf{w}}_{j}))+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))]$
@@
39:          $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow\max\left(0,\,\ell(\mathbf{w}_{j}^{(s)})-\ell(\hat{\mathbf{w}}_{j})-\epsilon\right)$
40:          $g_{\bar{\mathcal{P}}}^{t}\leftarrow g_{\bar{\mathcal{P}}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
41:      end if
42:      $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\alpha_{\mathcal{P}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\mathcal{P},scaled}^{(s)})$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\alpha_{\bar{\mathcal{P}}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\bar{\mathcal{P}},scaled}^{(s)})$triangleright Mirror ascent
43:      $Z\leftarrow\alpha_{\mathcal{P}}^{(s+1)}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}$
44:      $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\frac{\alpha_{\mathcal{P}}^{(s+1)}}{Z}$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\frac{\alpha_{\bar{\mathcal{P}}}^{(s+1)}}{Z}$triangleright Weight normalization
@@
45:      $\mathbf{w}_{t}^{(s+1)}\leftarrow\mathbf{w}_{t}^{(s)}-\eta_{\mathbf{w}}(\alpha_{\mathcal{P}}^{(s+1)}g_{\mathcal{P}}^{t}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}g_{\bar{\mathcal{P}}}^{t})$$\forall t\in\mathcal{T}$$\triangleright$ Mirror descent
46:end for
47:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$

### v3

@@ Appendix A Adapting FairPO for Multi-Attribute Generation
This section outlines our planned extension of FairPO to multi-attribute generation, a conceptual direction for future work. The goal is to generate a sequence $y$ from a prompt $x$ using a policy $\pi_{\mathbf{w}}(y|x)$ that aligns with fairness goals over a set of attributes $\mathcal{A}$. The core idea involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets and retaining the GRPO minimax structure (Eq. 18). The group losses would be defined over a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$, with preference losses like DPO applied to the log-probabilities of entire generated sequences rather than individual label scores.
Proposed Privileged Loss ($\mathcal{L}_{\mathcal{P}}$): For privileged attributes $j\in\mathcal{P}$, the goal is to ensure the learned policy $\pi_{\mathbf{w}}$ strongly reflects preferences $y_{w}\succ y_{l}$ established by that attribute. This is achieved using a standard DPO loss, averaged over the privileged subset of the preference data:
@@
Minimizing this loss directly encourages the model to favor preferred sequences for preferences driven by privileged attributes, relative to the reference policy $\pi_{\text{ref}}$.
Proposed Non-Privileged Loss ($\mathcal{L}_{\bar{\mathcal{P}}}$): For non-privileged attributes $k\in\bar{\mathcal{P}}$, the objective remains analogous to the classification setting: preventing significant performance degradation. This is accomplished with a hinge formulation based on the DPO loss:
@@
This penalizes the model only if its preference modeling for non-privileged attributes degrades substantially beyond baseline performance (represented by $\log 2$ for random preference) plus a slack $\epsilon^{\prime}$. The overall FairPO objective would then use GRPO to balance these two losses.
Planned Experiments: The planned task is generating text from a prompt, requiring preference datasets where the driving attribute $j_{i}$ is known. This could involve partitioning attributes like ‘Helpfulness’ or ‘Harmlessness’ from existing RLHF datasets (e.g., Anthropic HH-RLHF [Bai and Others, 2022], Summarization preferences [Stiennon et al., 2022]) or creating new, explicitly annotated data. We plan to fine-tune pretrained language models (e.g., Gemma [Team, 2024], Llama [Touvron and Others, 2023], Mistral [Jiang and Others, 2023]) with this objective. Evaluation would rely on preference-based metrics, such as win rates for privileged and non-privileged attribute groups, compared against baselines like standard DPO and SFT. Future experiments would also explore hyperparameter effects and adapting the SimPO/CPO loss formulations for generation.
@@ Appendix B FairPO Algorithm
The FairPO framework is trained iteratively to solve the minimax objective presented in Eq. 18. The detailed procedure, which is inspired by the DPO-based variant of FairPO, is provided in Algorithm  2.
@@
Initialization: The training process begins by initializing the model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$, for instance by copying them from a pre-trained reference model $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$. The adaptive group weights, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$, are typically set to uniform values such as 0.5 each.
Iterative Training Loop: The core of the framework is an iterative training loop. In each step, an instance $(x_{i},y_{i})$ is sampled from the dataset $\mathcal{D}$, and a single label $r\in\mathcal{T}$ is randomly selected from that instance for processing. The subsequent steps depend on whether this sampled label belongs to the privileged or non-privileged set.
If the sampled label $r$ is in the privileged set $\mathcal{P}$, the algorithm first identifies if a confusing set $S_{il}$ exists for that label (where $l=r$), as detailed in Algorithm 2. The loss computation is then conditional on this set:
  • If confusing examples exist ($S_{il}\neq\emptyset$), a DPO-inspired preference loss is computed between label $l$ and a randomly sampled confusing example $k\in S_{il}$. This preference loss directly encourages the model to improve its ranking of $l$ relative to its specific confounder $k$.
  • If no confusing examples are found ($S_{il}=\emptyset$), the algorithm reverts to a standard base classification loss (e.g., BCE, Eq. 10) for label $l$. This fallback is crucial as it ensures the model continues to receive a learning signal on easier instances, promoting stable training.
The loss calculated from either of these cases contributes to the current step’s privileged group loss, $\mathcal{L}_{\mathcal{P}}^{(s)}$. Conversely, if the sampled label $r$ belongs to the non-privileged set $\bar{\mathcal{P}}$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed according to Eq. 17. This loss penalizes the model only if its performance on the label $j=r$ deviates from the reference model’s performance by more than a predefined slack margin $\epsilon$.
@@
After the appropriate group loss is computed, the GRPO mechanism performs two key updates. First, the Adaptive Weight Update adjusts the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ using a mirror ascent step. This step uses an exponential weighting based on the current (and scaled) group losses, dynamically increasing the focus on the group that is currently performing worse (Lines 39-41). Second, the Model Parameter Update updates all model parameters $\mathbf{w}_{t}$ via a mirror descent step, using a combined gradient that is weighted by the newly updated adaptive weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$.
@@
This entire process repeats for a predefined number of iterations or until convergence, allowing FairPO to dynamically balance its objectives to achieve robust fairness. For variants like FairPO-SimPO or FairPO-CPO, the core logic remains identical; only the DPO-inspired preference loss component is replaced with their respective preference formulations (e.g., Eq. 12 or 13). The overall GRPO structure and non-privileged handling are consistent across all variants.
@@
1:Initialize: $\{\mathbf{w}_{t}^{(0)}\in\mathbb{R}^{d}|\forall t\in\mathcal{T}\}$ (e.g., copy $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$), $\alpha_{\mathcal{P}}^{(0)}\leftarrow 0.5,\alpha_{\bar{\mathcal{P}}}^{(0)}\leftarrow 0.5$.
2:Choose: $\eta_{\mathbf{w}},\eta_{\alpha}$, $\beta$, $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$, $\epsilon$.
@@
10:     $l\leftarrow r$, $S_{il}^{\text{neg}}\leftarrow\emptyset$, $S_{il}^{\text{pos}}\leftarrow\emptyset$
11:     if $y_{il}=+1$ then $\triangleright$ True Positive case for privileged label $l$
12:      $S_{il}^{\text{neg}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\geq m(x_{i};\mathbf{w}_{l}^{(s)})\}$, $S_{il}\leftarrow S_{il}^{\text{neg}}$
@@
13:     else if $y_{il}=0$ then $\triangleright$ True Negative case for privileged label $l$
14:      $S_{il}^{\text{pos}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\leq m(x_{i};\mathbf{w}_{l}^{(s)})\}$, $S_{il}\leftarrow S_{il}^{\text{pos}}$
15:     end if
@@
16:     if $S_{il}\neq\emptyset$ then $\triangleright$ Confusing examples exist, use DPO-inspired loss
17:      Sample $k\in S_{il}\sim Uniform(\frac{1}{|S_{il}|})$
18:      if $y_{il}=+1$ then $\triangleright$ DPO for True Positive $l$ vs Confusing Negative $k$
19:        $h_{\mathbf{w}^{(s)}}(x_{i},l,k)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)$.
@@
20:        $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},l,k)\right)$
21:      else if $y_{il}=0$ then $\triangleright$ DPO for True Negative $l$ vs Confusing Positive $k$
22:        $h_{\mathbf{w}^{(s)}}(x_{i},k,l)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)$.
@@
23:        $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},k,l)\right)$
24:      end if
@@
25:      $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{pref}}$, $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{pref}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
26:     else $\triangleright$ No confusing examples, use BCE loss for privileged label $l$
27:      $\mathcal{L}_{\text{BCE}}\leftarrow-[y_{il}\log m(x_{i};\mathbf{w}_{l}^{(s)})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}^{(s)}))]$
28:      $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{BCE}}$, $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{BCE}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
29:     end if
30:    else if $r\in\bar{\mathcal{P}}$ then $\triangleright$ Handle non-privileged label
31:     $j\leftarrow r$
32:     $\ell(\mathbf{w}_{j}^{(s)})\leftarrow-[y_{ij}\log(m(x_{i};\mathbf{w}_{j}^{(s)}))+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}^{(s)}))]$
33:     $\ell(\hat{\mathbf{w}}_{j})\leftarrow-[y_{ij}\log(m(x_{i};\hat{\mathbf{w}}_{j}))+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))]$
@@
34:     $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow\max\left(0,\,\ell(\mathbf{w}_{j}^{(s)})-\ell(\hat{\mathbf{w}}_{j})-\epsilon\right)$, $g_{\bar{\mathcal{P}}}^{t}\leftarrow g_{\bar{\mathcal{P}}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
35:    end if
36:    $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\alpha_{\mathcal{P}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\mathcal{P},scaled}^{(s)})$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\alpha_{\bar{\mathcal{P}}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\bar{\mathcal{P}},scaled}^{(s)})$ $\triangleright$ Mirror ascent
37:    $Z\leftarrow\alpha_{\mathcal{P}}^{(s+1)}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}$, $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\frac{\alpha_{\mathcal{P}}^{(s+1)}}{Z}$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\frac{\alpha_{\bar{\mathcal{P}}}^{(s+1)}}{Z}$ $\triangleright$ Weight normalization
@@
38:    $\mathbf{w}_{t}^{(s+1)}\leftarrow\mathbf{w}_{t}^{(s)}-\eta_{\mathbf{w}}(\alpha_{\mathcal{P}}^{(s+1)}g_{\mathcal{P}}^{t}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}g_{\bar{\mathcal{P}}}^{t})$$\forall t\in\mathcal{T}\triangleright$ Mirror descent
39:end for
40:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$